Transformation in graph of quadratic equation

Objective: to analyzed the effect of changing the values of a,h,and k of a quadratic function and it's graph

Direction: to determine the transformation in the graph of an equation analyzed the effect of changing the values of a,h,and k in the graph of y=x²


1.y=5(x+9)²+11
2.y=-7(x-10)²+4
1
Expert's answer
2020-12-08T09:57:44-0500

"y=5(x+9)^2+11"

1. "y_1=x^2" black parabola

2. Shift the graph of "y_1=x^2" to the left 9 units

"y_2=(x+9)^2" blue parabola.

3. Stretch the graph "y_2=(x+9)^2" vertically by factor of 5

"y_3=5(x+9)^2" green parabola.

4. Shift the graph of "y_3=5(x+9)^2" upward 11 units

"y_4=5(x+9)^2+11" red parabola.





"y=-7(x-10)^2+4"

1. "y_1=x^2" black parabola

2. Shift the graph of "y_1=x^2" to the right 10 units

"y_2=(x-10)^2" blue parabola.

3. Stretch the graph "y_2=(x-10)^2" vertically by factor of 7

"y_3=7(x-10)^2" green parabola.

4. Reflect the graph of "y_3=7(x-10)^2" in the x-axis

"y_4=-7(x-10)^2" orange parabola.

4. Shift the graph of "y_4=-7(x-10)^2" upward 4 units

"y_5=-7(x-10)^2+4" red parabola.